OpenStudy (anonymous):

principal value of the argument -7/61+16/61(i) in the range of -pi to pi help anyone?

OpenStudy (anonymous):

solve \[\tan(\theta)=\frac{b}{a}\]

OpenStudy (anonymous):

\[a=\frac{-7}{61},b=\frac{16}{61}\] \[\frac{b}{a}=-\frac{16}{7}\]so \[\theta=\tan^{-1}(-\frac{16}{7})\]

OpenStudy (anonymous):

now you need a calculator

OpenStudy (anonymous):

in the answer it says pi - tan^-1 (16/7)

OpenStudy (anonymous):

that's what confused me

OpenStudy (anonymous):

well if you don't mind me saying that is a really really stupid answer that tells you nothing

OpenStudy (anonymous):

analysis IS STUPID i have an exam tomorrow haha

OpenStudy (anonymous):

just tells you that \[\pi-\tan^{-1}(\frac{16}{7})=\tan^{-1}(-\frac{16}{7})\]

OpenStudy (anonymous):

that's not the same is it?

OpenStudy (anonymous):

inverse tangent already has range \[[-\frac{\pi}{2},\frac{\pi}{2}]\]

OpenStudy (anonymous):

i am sticking with my answer.

OpenStudy (anonymous):

actually now that i look at it i don't even understand the question "has range -pi to pi"

OpenStudy (anonymous):

-pi to pi is the whole cirlce

OpenStudy (anonymous):

2 pi is a circle

OpenStudy (anonymous):

oooooooooooooooh i see you are in quadrant 2

OpenStudy (anonymous):

it says the principal value of the argument is in the range (-pi,pi]

OpenStudy (anonymous):

got it. ok fine

OpenStudy (anonymous):

you are in quadrant 2 so angle is between \[\frac{\pi}{2}\] and \[\pi\]

OpenStudy (anonymous):

add pi to \[\tan^{-1}(-\frac{16}{7})\]

OpenStudy (anonymous):

thanks, how would i know when to add pi?